You have a deck of 52 playing cards.

(i) How many different 8 card hands can be dealt?

(ii) What is the probability that a hand of 8 dealt randomly contains (exactly) 2 aces?

(iii) What is the probability that a hand of 7 dealt randomly will have 7 cards of the same

suit?


An influenza vaccine is produced by two different companies. It is known that a vaccine

produced by company 1 is effective with probability 0:89, while a vaccine produced by

company 2 is effective with probability 0:93. We also know that company 1 supplies 40%

of the vaccines, while company 2 supplies 60% of the vaccines ordered by the government.

(i) What is the probability that a vaccine is effective, given that it was produced by company2?

(ii) What is the probability that a randomly chosen vaccine from the government’s order

is not effective?

(iii) What is the probability that given a vaccine is not effective that it was produced by

company 1?



1
Expert's answer
2021-12-12T18:03:51-0500

"a)"

"i)"

We apply combinations to find the 8 card hands can be dealt as follows,

number of ways ="\binom{n}{k}={n!\over(k!(n-k)!)}"

From the question, "n=52, k=8"

Therefore,

"\binom{52}{8}={52!\over(8!\times44!)}=752538194"

Thus, there are 752538194-8card hands that can be dealt.


"ii)"

To get exactly 2 aces, we need to choose 2 of the 4 aces and 6 of the other 48 cards. The number of ways to do that is

"\binom{4}{2}\times\binom{48}{6}={4!\over(2!*2!)}*{48!\over(6!*42!)}=6\times12271512=73629072"

Therefore, the probability is, "{73629072\over752538194}=0.0978(4dp)"


"iii)"

We first choose 7 cards out of a total of 13 cards in a suit. But we only want one out of 4 suits. Therefore, the number of ways for choosing 7 cards out of a total of 13 cards is "\binom{13}{7}={13!\over(7!\times6!)}=1716" and the number of ways of choosing 1 out out 4 suits is, "\binom{4}{1}={4!\over(1!\times3!)}=4"

Total number of 1716*4=6864

Hence the probability is, "{6864\over752538194}=9.121132e^{-6}"

Therefore, the probability  that a hand of 7 dealt randomly will have 7 cards of the same

suit is =9.121132e^{-6}


"b)"

Let "C_1" and "C_2" be the events that the influenza vaccine is manufactured by company 1 and 2 respectively. Also, let "E" be the event that the influenza vaccine produced is effective. "E'" is the event that the influenza vaccine produced is not effective.

The following probabilities are given,

"p(E|C_1)=0.89,\space p(E|C_2)=0.93,\space p(C_1)=0.40,\space p(C_2)=0.60"


"i)"

The probability that a vaccine is effective, given that it was produced by company 2 is given as,

"p(E|C_2)=0.93" as stated above.


"ii)"

We need to determine the probability that a randomly selected vaccine is effective. To do so, we shall apply the law of total probability as follows,

"p(E)=p(E|C_1)*p(C_1)+p(E|C_2)*p(C_2)=0.89*0.40+0.93*0.60=0.356+0.558=0.914"

The probability that a randomly selected vaccine is not effective is given as,

"p(E')=1-p(E)=1-0.914=0.086"

Therefore, the probability that a randomly selected vaccine is not effective is 0.086.


"iii)"

We determine the conditional probability, "p(C_1|E')" defined as,

"p(C_1|E')={p(C_1\cap E')\over p(E')}"

"P(C_1\cap E')=0.4\times0.11=0.044\space and \space p(E')=0.086"

Thus, "p(C_1|E')={0.044\over 0.086}=0.5116(4dp)"

Therefore, the probability that a randomly selected vaccine is produced by company 1 given it is not effective is 0.5116.